The manager of a restaurant in a large city claims that waiters working in all restaurants in his city earn an average of $150 or more in tips per week. A random sample of 25 waiters selected from restaurants of this city yielded a mean of $139 in tips per week with a standard deviation of $28. Assume that the weekly tips for all waiters in this city have a normal distribution.





1
Expert's answer
2021-12-16T16:17:56-0500

Null hypothesis "H_0:\mu=150."

Alternative hypothesis "H_a:\mu<150."

Test statistic: "t=\frac{\bar x-\mu}{\frac{s}{\sqrt{n}}}=\frac{139-150}{\frac{28}{\sqrt{25}}}=-1.96."

Degrees of freedom: "df=n-1=25-1=24."

P-value: "p=P(T<-1.96)=0.0309."

Since the p-value is less than 0.05, reject the null hypothesis.

There is a sufficient evidence that waiters working in all restaurants in his city earn an average less than $150.